# Benchmarking Differentially Private Tabular Data Synthesis

### v1

@@ 8. Experimental Setup
Our evaluations aim to (1) compare all well-established methods within our unified framework; (2) explore and verify the importance of preprocessing in DP tabular synthesis tasks; (3) investigate feature selection and synthesis modules of these methods for a more fine-grained comparison. The comparison results can validate previous theoretical analysis and guide method selection for different modules within the framework.
To achieve this, we design three main experiments as follows, (1) Overall Evaluation: We evaluate method performance across metrics and datasets; (2) Preprocessing Investigation: These experiments focus on the preprocessing investigation by comparing preprocessed datasets with raw ones; (3) Module Comparison: We evaluate the effectiveness of different feature selection and synthesis modules by reconstructing algorithms using them.
@@ 8.1. Datasets
The selection of datasets is crucial in comparison. To make a comprehensive comparison, we would expect that (1) the datasets should vary in record size, attributes’ domain sizes, and attribute distribution to better demonstrate the capacity of methods in synthesizing datasets of different complexity; (2) the proportion of numerical attributes and categorical attributes should be different to evaluate different preprocessing methods better. Therefore, we choose five datasets used in previous work (Liu et al., 2023; McKenna et al., 2022; Zhang et al., 2021; Kotelnikov et al., 2023) as our datasets, which are ACSincome (INC), ACSemploy (EMP), Bank (BK), Higgs-small (HIG) and Loan (LN). The detailed information about these datasets is provided in the appendix.
@@
ML Efficiency $\uparrow$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
PrivSyn $0.38$ $0.39$ $0.41$ $0.42$ $0.43$ $0.39$ $0.47$ $0.47$ $0.47$ $0.40$ $0.43$ $0.43$ $0.25$ $0.26$ $0.26$
@@
Ground Truth $0.05$ $0.05$ $0.05$ $0.02$ $0.02$ $0.02$ $0.03$ $0.03$ $0.03$ $0.12$ $0.12$ $0.12$ $0.10$ $0.10$ $0.10$
8.2. Implementations
In our experiments, we consider PrivSyn, PrivMRF, RAP++, AIM, Private-GSD, GEM, DP-MERF, and TabDDPM. We do not include RAP in the experiments because RAP++ directly improves upon it. Moreover, notice that Private-GSD is a synthesis algorithm, we use its one-shot 2-way marginal version in the overall evaluation, which is also used in their original work, and pay more attention to its performance in module comparison. Finally, we train TabDDPM under DP-GSD by opacus (Yousefpour et al., 2021). We repeat all evaluations five times and report the average results. The DP parameter $\delta$ is set to be $10^{-5}$ by default. PrivSyn and AIM are executed on CPUs, while the other methods are executed with GPUs. We by default employ uniform binning and rare category merging preprocessing methods. The detailed hyperparameters setting of studied methods, including preprocessing algorithms, is provided in the appendix.
8.3. Evaluation Metrics
Various evaluation approaches have been proposed (Zhang et al., 2021; McKenna et al., 2022; Du and Li, 2024). However, most current works focus on comparing synthesis utility, or how similar synthetic data is to real data. While utility is important, algorithm efficiency is also a critical factor in algorithm selection. Therefore, we mainly consider the following metrics.
Machine Learning Efficiency (higher is better). A widely accepted metric for evaluating generated data is its performance on downstream tasks. A common approach involves training Machine Learning (ML) models on the generated data and assessing their performance on test data.
The previous works (Zhang et al., 2021; Kotelnikov et al., 2023; Du and Li, 2024) typically select one or more ML models as downstream tasks. It’s important to note that a large number of models do not necessarily lead to a fairer evaluation. Generally, simpler models have weaker data-fitting capabilities, which cannot reach fair conclusions. In our comparison, we therefore selected four machine learning models known for strong performance across various datasets: MLP, CatBoost, XGBoost, and Random Forest. We report the average F1 score on held-out test data as our metric value, and other related metrics, such as AUC and accuracy, are provided in the appendix.
Query Error (lower is better). Making queries (Chatfield, 2018) is a commonly used data analysis technique, which can also be conducted to measure relatively high-dimensional similarity due to its high efficiency. Here, we consider using the 3-way marginal query method employed by Du et al. and Mckenna et al. (Du and Li, 2024; McKenna et al., 2019), which utilizes the statistical $\ell_{1}$ error of frequency query result to reflect the magnitude of the error. Formally, the query error can be expressed as, $\mathbb{E}_{r\in R}\left|q_{r}(D_{syn})-q_{r}(D_{test})\right|,$ where $q_{r}$ refers to the query function, which is a combination of range query (for numerical attributes) and point query (for categorical attributes). $\mathbb{E}$ is the mathematical expectation, and $R$ refers to the set of all 3-way marginals.
Fidelity Error (lower is better). Marginal Fidelity is precise in evaluating low-dimensional similarity, such as average 2-way marginal discrepancy. For example, the work by Du et al. (Du and Li, 2024) suggests using Wasserstein distance as the fidelity measurement, but this method can be infeasible in terms of computation time when dealing with complex data. Alternatively, total variation distance (TVD) can be used for measurement, which has also been utilized in some studies (Tao et al., 2022; Tran and Xiong, 2024). We define the TVD as $\frac{1}{2}\sum_{1\leq i\leq j\leq d}\left|M^{\text{syn}}_{i,j}-M^{\text{test}}_{i,j}\right|,$ where $M^{\text{syn}}_{i,j}$ and $M^{\text{test}}_{i,j}$ are the real 2-way marginals determined by the synthetic dataset and test dataset, respectively.
In addition to our primary evaluation criteria, we also use some secondary metrics, such as marginal size. These metrics are simple enough, so we omit the discussion of them here.
9. Experimental Results
9.1. Overall Evaluation
Utility Comparison. Table 2 shows the detailed utility metrics of different methods respectively. To make the result more straightforward, we utilize the t-SNE (Van der Maaten and Hinton, 2008), a dimensionality reduction technique, to visualize the synthesis results of all methods on the Bank dataset in Figure 2. By reducing the dimensionality of the dataset with t-SNE technique and plotting the scatter distribution, these visualizations reveal the distribution overlap between synthetic and real data, demonstrating synthesizers’ ability to generate data resembling the original.
The first clear conclusion is that PrivMRF and AIM achieve the best utility metrics among all methods. Another straightforward finding is that statistical methods generally perform better than machine learning approaches. Both AIM and PrivMRF rely on graphical models, and the quantitative evaluations of GEM, DP-MERF, and TabDDPM are worse than most other statistical methods in some metrics, such as RAP++ and PrivSyn in fidelity error.
@@ 9.1. Overall Evaluation
Figure 3. Scaled average utility and efficiency of different algorithms. Utility is obtained by the average performance of three utility metrics and efficiency is measured by average logarithmic running time. All measurements are scaled to $[0,1]$, and the higher value means better performance.
Utility & Efficiency Trade-off. We present the relationship between utility and efficiency on different algorithms in Figure 3. Although AIM and PrivMRF exhibit excellent utility performance, they require more execution time, representing a trade-off in utility. PrivMRF is slightly more efficient than AIM because it involves initialized marginal selection, which reduces the number of selection rounds and accelerates the algorithm. Additionally, all machine learning-based methods, despite having lower generation utility, demonstrate high efficiency, partly due to the use of GPUs in the model-fitting steps.
TabDDPM has a weak performance in synthesis utility. Among all the methods, TabDDPM performs poorly in all three dimensions, and in Figure 2, the visualization of the synthesis result does not match the real data distribution. We track the training loss with DP-SGD and compare it with the loss without DP-SGD. In Figure 4(a), we can observe that the loss under DP-SGD does not converge well when training. This raises our suspicion of DP-SGD’s efficiency. However, the diffusion model still demonstrates excellent performance when generating images (Li et al., 2024; Dockhorn et al., 2023) under DP-SGD. Therefore, we believe the failure of TabDDPM is potentially caused by its unsuitability for tabular data synthesis due to tabular data’s high sensitivity to exact attribute values and the relation between records (e.g., overall distribution), which is hard to achieve under DP-SGD.
(a) Training loss of TabDDPM with and without DP on Bank dataset under $\varepsilon=1.0$
(b) TVD between generated data and real data when training DP-MERF
(c) TVD between generated data and real data when training GEM
@@
Figure 5. Metrics under different preprocessing settings. These metrics are obtained by comparing preprocessed raw data with test data. Numerical preprocess means only conducting numerical discretization, and categorical preprocess means only conducting categorical rare merging. By default, the results are under the setting that $\varepsilon=1.0$ and $10\%$ of the budget is used for preprocessing (if needed).
@@
PrivSyn works poorly on machine learning efficiency. Another finding is that PrivSyn performs well on range query error and fidelity error metrics, but underperforms in terms of machine learning efficiency. This scenario aligns with our previous analysis in Section 7, which expresses the concern that GUM focuses more on local marginal cliques and may ignore global relevance. The experiments on reconstructed algorithms in Section 9.3 can also support this statement.
@@ 9.2. Preprocessing Investigation
Influence of Preprocessing on Dataset Information. Directly comparing the algorithms’ performances with and without preprocessing is difficult because running algorithms on raw datasets is often too time-consuming and computationally complex (e.g., some methods even require more than 24 hours to run on raw datasets). Therefore, we consider comparing the marginal sizes and utility metrics calculated on preprocessed and raw datasets. The detailed results are shown in Figure 5.
A straightforward finding is that preprocessing decreases the complexity of data, with the introduction of only a small error. In Figure 5, the average marginal size significantly decreases after preprocessing (from $10^{8}$ to $10^{3}$ in Higgs-small and Loan datasets). Meanwhile, its negative influences on utility are small enough (change on query error $<$ 0.003, TVD $<$ 0.1). We infer this is because binning can preserve most numerical characteristics, and the low-frequency categorical values only contribute a small proportion of the overall correlation between attributes in the dataset.
@@
Higgs-small $4870$ $6$ $100$ $60696$ $18$ $100$
Loan $101$ $9$ $100$ $93995$ $32$ $100$
Numerical Discretization Ablation. Firstly, we summarize the domain sizes under different discretization results in Table 3. It is obvious that PrivTree always generates significantly fewer bins to represent numerical attributes, even for attributes with highly complex value distributions. This approach not only reduces noise in feature measurements but also simplifies the features, raising concerns about losing information. To better demonstrate the influence of discretization, we conduct the ablation study on Higgs-small dataset, which contains the most complex numerical variables among all datasets. For this ablation, we focus on $\varepsilon=1.0$, as is standard practice (Du and Li, 2024; Tao et al., 2022). The results are shown in Table 4.
In most cases, PrivTree performs better on machine learning tasks but is slightly worse in query errors and fidelity errors compared with uniform binning. We believe fewer bins help capture overall correlations by introducing less noise, which benefits machine learning tasks. However, query tasks and fidelity are more sensitive to exact values, and more bins lead to higher accuracy. Two exceptions are Private-GSD and GEM, where PrivTree outperforms uniform binning. We hypothesize that this is because PrivTree reduces the dimensionality of variables, thereby facilitating network training and improving the convergence of the genetic algorithm.
These results do not completely align with observations in Tao et al.’s work (Tao et al., 2022). We would propose three possible reasons. Firstly, Higgs-small dataset is more value-rich than the datasets used in their work, which can lead to high sensitivity to binning methods. Secondly, the evaluation metrics are different, while ours focus more on measurements on higher-dimensional marginals (e.g., 3-way marginals). This can influence the results. Finally, the investigated algorithms vary in the two works, which may lead to different comparison conclusions. Another unusual finding is that under PrivTree binning, AIM shows a significantly worse runtime. We believe this is because smaller domain sizes will cause AIM to allocate budget to more marginals. This may lead to large cliques, slowing down the execution speed of PGM.
Table 4. Marginal-based methods’ performance based on different discretization methods. The result is obtained on the Higgs-small dataset and under $\varepsilon=1.0$.
Method ML Efficiency $\uparrow$ Query Error $\downarrow$ Fidelity Error $\downarrow$ Running Time $\downarrow$
PrivTree Uniform PrivTree Uniform PrivTree Uniform PrivTree Uniform
PrivSyn $\mathbb{0.43}$ $\mathbb{0.43}$ $0.005$ $\mathbb{0.004}$ $\mathbb{0.19}$ $0.20$ $\mathbb{2}$ min $6$ min
PrivMRF $\mathbb{0.65}$ $0.64$ $0.005$ $\mathbb{0.003}$ $0.19$ $\mathbb{0.16}$ $15$ min $\mathbb{7}$ min
AIM $\mathbb{0.67}$ $0.65$ $0.005$ $\mathbb{0.003}$ $0.20$ $\mathbb{0.19}$ $690$ min $\mathbb{18}$ min
RAP++ $\mathbb{0.55}$ $0.53$ $0.030$ $\mathbb{0.029}$ $\mathbb{0.55}$ $0.65$ $71$ min $\mathbb{37}$ min
Private-GSD $\mathbb{0.50}$ $0.47$ $0.043$ $\mathbb{0.044}$ $\mathbb{0.57}$ $0.65$ $\mathbb{57}$ min $\mathbb{57}$ min
GEM $\mathbb{0.56}$ $0.54$ $\mathbb{0.019}$ $0.061$ $\mathbb{0.29}$ $0.55$ $\mathbb{0.4}$ min $6$ min

### v2

@@ 8. Experimental Setup
Our evaluations aim to (1) compare all well-established methods within our unified framework; (2) explore and verify the importance of preprocessing in DP tabular synthesis tasks; (3) investigate feature selection and synthesis modules of these methods for a more fine-grained comparison. The comparison results can validate previous theoretical analysis and guide method selection for different modules within the framework. To achieve these goals, we design and conduct three main experiments. We summarize them and some important findings as follows.
  1. Overall Evaluation: We evaluate method performance across metrics and datasets. A key observation is that statistical methods like AIM and PrivMRF demonstrate better synthesis utility, while under current implementations, deep learning methods can achieve higher time efficiency.
  2. Preprocessing Investigation: These experiments focus on the preprocessing investigation by comparing preprocessed datasets with raw ones. Through experiments, we prove that preprocessing is essential in reducing algorithm complexity without introducing much synthesis error.
  3. Module Comparison: Finally, we evaluate the effectiveness of different feature selection and synthesis modules by fixing one module and reconstructing algorithms by changing the other module. We find that (1) the adaptive selection strategy can help the algorithm achieve a higher synthesis utility; (2) existing different synthesis modules show their limitations in either utility or scalability.
@@ 8.1. Datasets
To make a comprehensive comparison, we would expect datasets to (1) vary in record and domain size, and attribute distribution to better show method performance on datasets of various complexity, and (2) differ in the proportion of numerical and categorical attributes to assess the preprocessing methods. Therefore, we choose five datasets used in previous work (liu2023generating; mckenna2022aim; zhang2021privsyn; kotelnikov2023tabddpm) as our datasets, which are ACSincome, ACSemploy, Bank, Higgs-small and Loan. The brief information about these datasets is provided in Table 2.
Furthermore, in Figure 2, we plot the heat maps on absolute values of pairwise correlations across different datasets. Darker colors in the heat map mean higher correlation. We also plot the histograms of the attribute distribution’s Shannon entropy, defined as $H(X)\;=\;-\sum_{i=1}^{k}p_{i}\,\ln p_{i}$. The larger the entropy, the more informative and complex the distribution is. From Figure 2, we can observe that ACSincome and ACSemploy have stronger correlations between attributes, while attribute distributions in the other three datasets have higher entropy, indicating greater complexity.
Figure 2. Heat map of pairwise absolute correlations and histogram of attribute distribution’s Shannon entropy.
8.2. Implementations
In our experiments, we consider PrivSyn, PrivMRF, RAP++, AIM, Private-GSD, GEM, DP-MERF, and TabDDPM. We do not include RAP in the experiments because RAP++ directly improves upon it. Moreover, notice that Private-GSD is a synthesis algorithm; we use its one-shot 2-way marginal version in the overall evaluation, which is also used in their original work, and pay more attention to its performance in module comparison. Finally, we train TabDDPM under DP-GSD by Opacus (opacus). We repeat all evaluations five times and report the average results. The DP parameter $\delta$ is set to be $10^{-5}$ by. We, by default, employ uniform binning and rare category merging preprocessing methods. The detailed hyperparameter settings of the studied methods, including preprocessing algorithms, are provided in the appendix of our full paper (chen2025benchmarking).
We keep our implementation of these methods aligned with their released versions. PrivSyn and AIM are executed on CPU, while the other methods are executed with GPU. This is because PrivSyn and AIM are designed for CPU and executing them on GPU may not be able to bring better performance. For instance, PrivSyn adjusts data in cliques one by one, which does not require high-dimensional computation. We also try AIM (PGM) on GPU (also with the officially released code from its authors) and compare the results with those from CPU. A brief result is shown in Table 3. We can observe that there is no significant improvement in algorithm performance after executing on GPU.
8.3. Evaluation Metrics
Various evaluation approaches have been proposed (zhang2021privsyn; mckenna2022aim; du2024towards). However, most current works focus on comparing synthesis utility, or how similar synthetic data is to real data. While utility is important, algorithm efficiency is also a critical factor in algorithm selection. Therefore, we mainly consider the following metrics.
Machine learning efficacy (higher is better). A widely accepted metric for evaluating generated data is training downstream machine learning (ML) models on the generated data and assessing their performance on test data. The previous works (zhang2021privsyn; kotelnikov2023tabddpm; du2024towards) typically select one or more ML models as downstream tasks. It’s important to note that a large number of models do not necessarily lead to a fairer evaluation. Generally, simpler models have weaker data-fitting capabilities, which cannot reach fair conclusions. In our comparison, we therefore selected four deep learning models known for strong performance across various datasets: MLP, CatBoost, XGBoost, and Random Forest. We report the average F1 score on held-out test data as our metric value, and other related metrics, such as AUC and accuracy, are provided in the appendix of our full paper (chen2025benchmarking).
Table 3. Results of AIM (PGM) on different hardware (CPU and GPU). These results are obtained on ACSincome dataset.
[…]
0.2 AIM (CPU) 0.77 0.0016 0.093 2.0 min
AIM (GPU) 0.75 0.0024 0.105 2.5 min
1.0 AIM (CPU) 0.78 0.0011 0.064 6.3 min
AIM (GPU) 0.78 0.0011 0.064 37.0 min
Query Error (lower is better). Making queries (chatfield2018introduction) is a commonly used data analysis technique, which can also be conducted to measure relatively high-dimensional similarity due to its high efficiency. Here, we consider using the 3-way marginal query method employed by Du et al. and McKenna et al. (du2024towards; mckenna2019graphical), which utilizes the statistical $\ell_{1}$ error of frequency query results to reflect the magnitude of the error. Formally, the query error can be expressed as, $\mathbb{E}_{r\in R}\left|q_{r}(D_{syn})-q_{r}(D_{test})\right|,$ where $q_{r}$ refers to the query function, which is a combination of range query (for numerical attributes) and point query (for categorical attributes). $\mathbb{E}$ is the mathematical expectation, and $R$ refers to the set of all 3-way marginals. Moreover, due to complex real applications on synthetic data, we also worry that simple frequency queries, even though fundamental, may be insufficient to cover them. Therefore, we also consider Conditional Query Error to simulate Join and GroupBy in SQL (fix one attribute as the key column and query the frequency of the other attribute). This metric serves as a supplementary evaluation reference, providing comprehensive comparisons.
Fidelity Error (lower is better). Marginal Fidelity is precise in evaluating low-dimensional similarity, such as average 2-way marginal discrepancy. Total variation distance (TVD) can be used for such measurement, which has also been utilized in some studies (tao2022benchmarkingdifferentiallyprivatesynthetic; tran2024differentially). We define the TVD as $\frac{1}{2}\sum_{1\leq i\leq j\leq d}\left|M^{\text{syn}}_{i,j}-M^{\text{test}}_{i,j}\right|,$ where $M^{\text{syn}}_{i,j}$ and $M^{\text{test}}_{i,j}$ are the real 2-way marginals determined by the synthetic dataset and test dataset, respectively.
In addition to our primary evaluation criteria, there are many other widely applied metrics for tabular data, such as Wasserstein distance (peyre2019computational) and Cramer’s V measure (cramer1999mathematical). We do not include these metrics in our evaluation mainly because of their feasibility in complex cases and overlap with existing metrics. For example, Wasserstein distance can be infeasible in terms of computation time when dealing with high-dimensional and complex data. Moreover, Cramer’s V measure and other correlation measurements cannot provide deterministic information about data distributions. In other words, even if two distributions are different, they can have similar correlation results. Furthermore, correlation information can be covered by fidelity error, which is based on a precise measurement of total distribution. We will further validate this in the appendix of our full paper (chen2025benchmarking).
@@ 8.3. Evaluation Metrics
ML Efficacy $\uparrow$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
PrivSyn $0.43$ $0.39$ $0.42$ $0.40$ $0.43$ $0.39$ $0.47$ $0.47$ $0.47$ $0.40$ $0.43$ $0.43$ $0.25$ $0.26$ $0.26$
@@
[…]
@@
9. Experimental Results
9.1. Overall Evaluation
Utility Comparison. We provide detailed results on machine learning efficacy, query error, and fidelity error in Table 4, and show some brief results of conditional query error in Table 5. To make the result more straightforward, we utilize the t-SNE (van2008visualizing), a dimensionality reduction technique, to visualize the synthesis results of all methods on the Bank dataset in Figure 3. By reducing the dimensionality of the dataset with t-SNE technique and plotting the scatter distribution, these visualizations reveal the distribution overlap between synthetic and real data, demonstrating synthesizers’ ability to generate data resembling the original.
The first clear conclusion is that PrivMRF and AIM achieve the best utility metrics among all methods. Another straightforward finding is that statistical methods generally perform better than deep learning approaches. Both AIM and PrivMRF rely on graphical models, and the quantitative evaluations of GEM, DP-MERF, and TabDDPM are worse than most other statistical methods in some metrics, such as RAP++ and PrivSyn in fidelity error.
Utility & Efficiency Trade-off. We present the relationship between utility and efficiency on different algorithms in Figure 5. Although AIM and PrivMRF exhibit excellent utility performance, they require more execution time, representing a trade-off in utility. PrivMRF is slightly more efficient than AIM because it involves an initialization step, reducing the number of selection rounds. Additionally, all deep learning-based methods, despite having lower generation utility, demonstrate high efficiency. This is partly because some statistical methods run on CPU and also because they involve large-scale computations.
TabDDPM has a weak performance in synthesis utility. Among all the methods, TabDDPM performs poorly in all three dimensions, and in Figure 3, the visualization of the synthesis result does not match the real data distribution. We track the training loss with DP-SGD and compare it with the loss without DP-SGD. In Figure 4(a), we can observe that the loss under DP-SGD does not converge well when training. This raises our suspicion of DP-SGD’s efficiency. However, the diffusion model still demonstrates excellent performance when generating images (li2024privimage; dockhorn2023differentiallyprivatediffusionmodels) under DP-SGD. Therefore, we hypothesize that the failure of TabDDPM is due to its incompatibility with tabular data under a DP-SGD framework. Unlike image data, which is characterized by dense and repetitive features (e.g., visual similarity among images of the same type), tabular data exhibits sparse and highly divergent characteristics (e.g., unique relationships between attributes), which present significant convergence challenges for DP-SGD.
(a) Training loss of TabDDPM with and without DP on Bank dataset under $\varepsilon=1.0$.
(b) TVD of marginals between generated data and real data when training DP-MERF.
(c) TVD of marginals between generated data and real data when training GEM.
@@ 9.1. Overall Evaluation
Table 5. Conditional query errors of different methods under $\varepsilon=1.0$.
Method ACSincome ACSemploy Bank Higgs-small Loan
Privsyn 0.019 0.035 0.033 0.029 0.058
PrivMRF 0.012 0.013 0.026 0.026 0.053
RAP++ 0.021 0.026 0.064 0.098 0.098
AIM 0.013 0.011 0.022 0.025 0.057
Private-GSD 0.016 0.014 0.039 0.110 0.093
GEM 0.026 0.028 0.053 0.069 0.077
DP-MERF 0.308 0.314 0.215 0.099 0.455
TabDDPM 0.092 0.072 0.136 0.109 0.143
Figure 5. Scaled utility and efficiency of different algorithms. Average utility is obtained by taking the average of ML efficacy, query error, and fidelity error after normalizing them to $[0,1]$, guaranteeing their equal contribution to the aggregated metric. Other metrics are directly obtained from the average value of normalized original metrics.
@@
PrivSyn works poorly on machine learning efficacy. Another finding is that PrivSyn performs well on range query error and fidelity error metrics, but underperforms in terms of machine learning efficacy. This scenario aligns with our previous analysis in Section 7, which expresses the concern that GUM focuses more on local marginal cliques and may ignore global relevance. The experiments in Section 9.3 can also support this statement.
Figure 6. Metrics under different preprocessing settings. These metrics are obtained by comparing preprocessed raw data with test data. Numerical preprocessing means only conducting numerical discretization, and categorical preprocessing means only conducting categorical rare merging. By default, the results are under the setting that total privacy budget $\varepsilon=1.0$.
@@ 9.2. Preprocessing Investigation
Influence of Preprocessing on Dataset Information. Directly comparing the algorithms’ performances with and without preprocessing is difficult because running algorithms on raw datasets is often too time-consuming and computationally complex (e.g., some methods even require more than 24 hours to run on raw datasets). Therefore, we consider comparing the marginal sizes and utility metrics calculated on preprocessed and raw datasets. The detailed results are shown in Figure 6. A straightforward finding is that preprocessing decreases the complexity of data, with the introduction of only a small error. In Figure 6, the average marginal size significantly decreases after preprocessing (from $10^{8}$ to $10^{3}$ in Higgs-small and Loan datasets). Meanwhile, its negative influences on utility are small enough (change on query error << 0.003, TVD << 0.1). We infer this is because binning can preserve most numerical characteristics, and the low-frequency categorical values only contribute a small proportion of the overall correlation between attributes in the dataset.
Table 6. Different algorithms’ performance on ACSincome with and without preprocessing under $\varepsilon=1.0$. Here, “Prep” means preprocessed data, and “Raw” means raw data.
Method ML Eff. $\uparrow$ Query Err. $\downarrow$ Fid Err. $\downarrow$ Time (min) $\downarrow$
Prep Raw Prep Raw Prep Raw Preps Raw
PrivSyn 0.39 0.39 0.002 0.002 0.12 0.12 0.2 0.2
PrivMRF 0.78 0.78 0.001 0.001 0.06 0.07 1.5 0.9
AIM 0.78 0.78 0.001 0.001 0.06 0.06 5.8 6.3
RAP++ 0.74 0.74 0.004 0.005 0.22 0.24 13.3 26.8
Private-GSD 0.77 0.77 0.002 0.003 0.20 0.21 9.7 24.9
GEM 0.70 0.68 0.014 0.017 0.25 0.28 0.1 0.1
DP-MERF 0.66 0.67 0.024 0.018 0.52 0.51 0.1 0.1
TabDDPM 0.41 0.40 0.063 0.068 0.76 0.80 6.7 4.1
Over-preprocessing Problem. In both Algorithm 1 and Algorithm 2, we apply a threshold $\beta$ to filter out those attributes with small domains to avoid “over-preprocessing”. In Table 6, we present the synthesis results for ACSincome with mandatory preprocessing, which does not need preprocessing under threshold $\beta$. We can observe that preprocessing does not significantly influence the algorithms’ utility. We believe that, because these datasets are inherently simple, preprocessing does not overly distort their information and thus has little impact on the synthesis results. A notable finding is that RAP++ and Private-GSD run faster on the preprocessed dataset. We hypothesize that this is because preprocessing combines similar or less informative values together, which allows these two methods to capture the data characteristics more easily and converge more quickly. In summary, preprocessing can be unnecessary for those simple datasets, but mandatory preprocessing is likely not to be detrimental to synthesis performance.
Numerical Discretization Ablation. Firstly, we summarize the domain sizes under different discretization results in Table 8. It is obvious that PrivTree always generates significantly fewer bins to represent numerical attributes, even for attributes with highly complex value distributions. This approach not only reduces noise in feature measurements but also simplifies the features, raising concerns about losing information. To better demonstrate the influence of discretization, we conduct the ablation study on Higgs-small dataset, which contains the most complex numerical variables among all datasets. For this ablation, we focus on $\varepsilon=1.0$, as is standard practice (du2024towards; tao2022benchmarkingdifferentiallyprivatesynthetic). The results are shown in Table 7.
Table 7. Marginal-based methods’ performance on the Higgs-small with different discretization under $\varepsilon=1.0$. Here, “Tree” means PrivTree and “Uniform” means uniform binning.
Method ML Eff. $\uparrow$ Query Err. $\downarrow$ Fid Err. $\downarrow$ Time (min) $\downarrow$
Tree Uniform Tree Uniform Tree Uniform Tree Uniform
PrivSyn $0.43$ $0.43$ $0.005$ $0.004$ $0.19$ $0.20$ $2.1$ $6.0$
PrivMRF $0.65$ $0.64$ $0.005$ $0.003$ $0.19$ $0.16$ $14.7$ $7.1$
AIM $0.67$ $0.65$ $0.005$ $0.003$ $0.20$ $0.19$ $689.6$ $18.5$
RAP++ $0.55$ $0.53$ $0.030$ $0.029$ $0.55$ $0.65$ $74.6$ $40.0$
Private-GSD $0.50$ $0.47$ $0.043$ $0.044$ $0.57$ $0.65$ $58.2$ $58.3$
GEM $0.56$ $0.54$ $0.019$ $0.061$ $0.29$ $0.55$ $0.4$ $6.1$
Figure 7. The maximum attribute domain size of Loan dataset under fixed merging thresholds and different $\varepsilon$. In the figure, $0.0\%$” means no fixed threshold (merge only by $3\sigma$ criteria), and “None” means raw data.
@@
Higgs-small $4870$ 66 $100$ $60696$ $18$ $100$
Loan $101$ 99 $100$ $93995$ $32$ $100$
In most cases, PrivTree performs better on deep learning tasks but is slightly worse in query errors and fidelity errors compared with uniform binning. We believe fewer bins help capture overall correlations by introducing less noise, which benefits deep learning tasks. However, query tasks and fidelity are more sensitive to exact values, and more bins lead to higher accuracy. Two exceptions are Private-GSD and GEM, where PrivTree outperforms uniform binning. We hypothesize that this is because PrivTree reduces the dimensionality of variables, thereby improving the convergence. These results do not completely align with observations in Tao et al.’s work (tao2022benchmarkingdifferentiallyprivatesynthetic). We would propose three possible reasons. Firstly, Higgs-small dataset is more value-rich than the datasets used in their work, which can lead to high sensitivity to binning methods. Secondly, the evaluation metrics are different, while ours focus more on measurements on higher-dimensional marginals (e.g., 3-way marginals). This can influence the results. Finally, the investigated algorithms vary in the two works, which may lead to different comparison conclusions. Another unusual finding is that under PrivTree binning, AIM shows a significantly worse runtime. We believe this is because smaller domain sizes will cause AIM to allocate budget to more marginals. This may lead to large cliques, slowing down the execution speed of PGM.
@@
Figure 8. Average scaled evaluations for different marginal selection modules (first row) and synthesis modules (second row) on different datasets. We use PGM as the common data synthesizer and PrivSyn as the common marginal selector. The running time is scaled by min-max linear normalization, and others are scaled using ground truth values provided in Table 4. The small datasets include ACSincome and ACSemploy, whereas the large datasets consist of Bank, Higgs-small, and Loan.